Minutes of the 12th Committee Meeting of the Dublin University Rifle Club

Moderator Steven Watterson

Minutes taker Mark Dennehy

Present Steven Watterson (Captain), Benjamin Jones (Treasurer), Mark Dennehy (Secretary/Match Secretary), John Keeney (Ammo&Equipment Officer), Alan Smyth (Chief Range Officer), Deirdre Ninaber (Ladies’ Coordinator)

Absent(with apologies) Conor McDermottroe (Public Relations Officer),

Absent Jason Sibley,

Date 30 October 2001 1840 – 2140

Distribution DURC Committee

Overview of topics

1 Meeting formality

Steven proposed that we run the meeting very formally as we had a lot of items to get through and noone wanted an all-night session. This was agreed to by all.

2 Capital Grant Meeting

Steven scheduled a meeting to deal with the Capital Grant for Tuesday 6th October at 1830 on the DURC Range (relocation to be discussed once we all arrive).

3 Code of Conduct

Steven suggested that a formalised document was unnecessary as common sense should suffice. Ben disagreed, saying that formalisation was needed so we could discipline people

4 Pony Club Training

if needed. The question of how to do so would also be covered. Alan said it was difficult to articulate the rules of ettiquette, Mark said it was also time consuming and that a formal document acted to cover us as well as to serve as a reference document. John suggested that it be included in Club regulations instead. Ben said it should be more explicit than that. A vote was called for. Explicit document

For  Against  Abstention
2  4  0

Motion rejected Covered in club regulations

For  Against  Abstention
5  1  0

Motion carried

Steven asked who looked after the regulations, Mark replied that that was his job as Secretary and agreed to bring them to the next committee meeting. Task: Bring regulations to next meeting (Mark)

4 Pony Club Training

Steven noted that we had looked at DUEC and Westmeath seperately to this point. We do one afternoon with DUEC a week (with two weeks to go) and one weekend every 3 weeks with Westmeath. More ROs need to assist. Steven felt a rota should be drawn up and asked if volunteers or conscription should be used. Mark said that Westmeath and DUEC should be seperate things to volunteer on and that we couldn’t ask ROs to give up lectures. Alan asked why we were doing this training and what did we gain from it? Ben pointed out that we need to watch DUEC’s time as it upsets normal shooting by delaying the 1800 detail. Mark suggested shooting at target holders 10m from the 25yd firing line and thus integrating the DUEC shooters with regular times. John said that this took away from our own members and Mark pointed out that DUEC shooters were DURC members. Alan said that we don’t exist to shoot non-ISSF events, Mark retorted that this was completely wrong, that the club has never specified what events we shoot. Steven listed three alternatives and called for a vote: Take over a full evening for prone 10m shooting

For  Against  Abstention
0  6  0

Motion rejected Use an afternoon for shooting prone 10m

For  Against  Abstention
4  2  0

Motion carried

The option of dropping support for DUEC was not voted on. However, the afternoon was moved from Thursday (1500-1800) to Friday by vote:

5 Shooting Trousers

Use Friday as the DUEC training day

For  Against  Abstention
4  2  0

Motion carried

Steven is to contact DUEC and arrange for this change. Mark said he could not assist with DUEC on Friday or Westmeath on Saturday due to an NRPAI conference for the ISC. Task: Contact DUEC and arrange day change (Steven)

5 Shooting Trousers

Steven was going to bring this up in relation to the capital grant, but diverted it to the Tuesday meeting.

6 Squads

Steven is still waiting responses for times for training from Liam Murray, John Keeney, Jason Sibley, Adrian Costigan, David Gallagher and Shane Kelly.

7 Liam Murray

Ben reports that Liam has missed his last few RO duty details. John said he had sent a copy of the duty roster to Liam. Mark points out that this is the CRO’s job to sort out and Alan agreed to take care of it. Task: Speak to Liam about his duty roster slots. (Alan)

8 Treasurers Report

Ben reports that we have about £300 available for use, with more available soon when we turn in some more receipts. Mark is to get petrol receipts for this. Task: Get receipts for petrol used in UCD Open (Mark) Capital Grant applications must be in by November 20th. Mark pointed out that we should get ours in early due to a new DUCAC treasurer.

8.1 RO Credits

Ben reported that the RO Credits were climbing again, especially Alans and Johns. Mark suggested that we should implement the limits system Ben, himself and Steven had discussed earlier. There were questions about the legality of credits in the first place. Ben noted that non-payment was only one problem, some ROs were not noting down ammo used in the RO credits book. At this point the options for limiting were a hard limit beyond which credit would not be extended and a soft one, beyond which interest

8 Treasurers Report

was charged. John suggested a soft limit beyond which extra RO duty was imposed on a proportional scale. A vote was called for Hard Limit

For  Against  Abstention
0  6  0

Motion rejected Soft Limit with Interest

For  Against  Abstention
0  6  0

Motion rejected Soft Limit with extra RO Duty

For  Against  Abstention
0  6  0

Motion rejected

At this point it was evident more discussion was needed!

Alan pointed out that extra RO Duty was difficult to arrange due to inflexibility of academic timetables. Ben said we shouldn’t really be using RO Credits except as a last resort anyhow, but Alan felt that was removing an RO privilege. Steven said that enforcement of sanctions was a problem. Mark suggested a “Name&Shame” policy. Steven asked if there were other options apart from Name&Shame and free slots in the duty roster to be filled by those in debt. Some options were discussed and dismissed. Mark restated that the root of the problem was that we could not enforce sanctions meaningfully. Alan suggested getting payment on credits every fortnight - Mark said that simply brought up the enforcement problem every fortnight. Mark said that John’s idea of extra RO Duty was the only solution that didn’t need enforcement of sanctions. John as the person owing the most was asked if that was acceptable, and he agreed to it. Soft limit with extra RO Duty

For  Against  Abstention
4 (3.5!)  0 (0.5!)  1

Motion carried (John voted half for and half against in jest!)

Amounts were discussed. £20 as a limit was considered and one duty night for every £10 over that limit. Ben pointed out that this would mean people would just accept £20 as the limit and not pay past that point. At 14 range officers, this was too much for the club to bear. Discussion on amounts followed, and it was finally agreed that Ben would post warning to ROs that they were over the limit and that if they did not pay by a set date, that they would owe one extra detail if they owed more than £5 up to £20; and one detail for every £20 over that that they owed :

Amount Owed Extra Details
£5 – £20 1
£20 – £40 2

and so on The time period for this was discussed, but not agreed on. It was suggested that this

was an operational point that could be fine-tuned later.

9 CROs Report

8.2 Noncapital grant

Ben reported the next application deadline was next week and wondered if anything was needed. Mark suggested upgrading estimates for entry fees.

9 CROs Report

The question of adding Eric to the photo list as an occasional RO was discussed and agreed. However, Eric had requested a key to the lounge. This was akward as the college only allows for Officers to have this key, and at present Steven has none. Consensus was to turn down the request. Give Eric a key

For  Against  Abstention
0  6  0

Motion rejected

10 Secretary’s Report

Mark reported that we had had no pertinant mail, and that we would be using the rifle@maths.tcd.ie email list due to its superior capabilities (archives, web-based admin, digests, etc)

11 A&E Report

John reported that we would soon require more Lapua Master (1000 rounds or so). We have lots of T22 and pellets. Air Rifle S is under repair with no time or money estimate available. Dave Cooney has reported he can test our sights. Alan pointed out he’d made a bad job of it before. Steven and John disagreed over whether the problem could be traced to the sights alone or head position. John insisted the sights were in dire need of servicing. Mark said Anschutz had a lifetime guarantee, though Steven wasn’t sure of this. Alan said that Mech. Eng. had equipment to allow for testing of the sights. It was agreed Alan and John would test them before we go to the trouble of sending them off for repair. Task: Test the sights in MechEng (John, Alan)

Targets came up as a topic. The NSRA ones we now have for air as 5-strip targets are fine for club training but Mark felt they would not suffice for competition and would lead to the same fiasco in the Open as happened in the Nationals’ Final. Ben said if we got Edelmann targets we had to get them locally due to prohibitive shipping costs. John replied that the previous high shipping costs was more probably due to the 50m and 25yd targets we’d ordered with them. Mark suggested trying the NSRA ones to find if they were unsuitable. John agreed to price Edelmann targets and Mark gave him the documentation they had sent him.

12 DURC 10m Open

The Priority zero purchasing list came up for discussion. John needed the bank draft to order the gear from the NSRA (for £128 sterling). Ben agreed to arrange a cheque

12 DURC 10m Open

The venue was agreed as UCDRC as WTSC was too expensive and difficult to get our people to. Steven was adamant that the event had to cover its own costs. It was agreed to be a 2-day event with prior shooting by arrangement. All ROs were to be available for duty and Steven felt they should be present throughout the weekend from enthuasism. A rota for RO duty should be drawn up. Scoring was mentioned - Keith will do one day, Steven volunteered for another, Ben mentioned data entry and Mark noted that we needed a minimum of one RO and two in the stats office. Reentrys would be allowed and a 10-shot final will be run. A new graduated guage was needed and added to the priority zero list for John to buy. It was agreed to discuss the fine detail at another meeting and the meeting for the Capital Grant was co-opted. Mark is going to try to find a sponsor for the event. Task: Find sponsorship for 10m Open (Mark) Task: Announce 10m Open (Mark)

13 RO Training Course

Mark reported that an NRPAI RO Training Course was being scheduled for November 17. The advantages of having nationally-qualified ROs was discussed, with mention made of the upcoming proposed risk assessment by DUCAC. It was decided to send six ROs on the course this time. Task: Allocate funds for six ROs to attend course (Ben)

14 RO Dinner

Mark asked if we could select a date for the RO Dinner. Steven asked what the format for the dinner was and that was explained. It was agreed to discuss the matter of the Committee dinner later. Mark gave a list of people we could expect to invite: Serving ROs, Past RO’s and committee, UCD’s Committee, the NTSA Committee and members of the National Squad. Alan said he felt it wrong to make a profit on the RO Dinner. Ben pointed out that this didn’t happen, that the extra cash usually went on the wine. A disussion ensued on the restaurant to select. Mark volunteered to acquire a selection of suitable menus, after which we could make a decision. This was agreed on. Task: Acquire menus for RO Dinner (Mark)

However a date still had to be selected. Balancing the timetables of students and alumni was discussed and eventually, after some lenghty debate, December 15 was selected. Steven was worried that we might be excluding our current members by choosing a late date.

15 RRPC vs DURC 40th Anniversary Shoot

The dinner for the 40th anniversary was also discussed in this vein. A lengthy debate on what “black tie” meant took place. It was decided however, that suits were all that were needed for the event, and that this would be reflected in the wording of the invitations. It was agreed to further discuss this next week.

15 RRPC vs DURC 40th Anniversary Shoot

Mark reported on the proposed anniversary shoot between RRPC and DURC. The calendar being allready full and our resources stretched, it was debated as to why this should be done at all. Alan felt that we should not do it as it was too great a strain on time and resources. John felt that it was a good idea to do it as it would strengthen relations with RRPC. Steven summarised that the pros were better relations, prior agreement with Nick West (RRPC) and that it looks good to DUCAC, and the cons were that we have a packed calendar and that RRPC and DURC aren’t demographically compatible. A vote was called : Hold the anniversary shoot

For  Against  Abstention
4  2  0

Motion carried

Mark is to liase with RRPC and select a date.

16 UCDRC/DURC/WTSC shoot

Ben brought up the three-club shoot that Geoff Cooney had been discussing last year, and wondered if the RRPC-DURC shoot would cause it to be cancelled. Mark pointed out that a solution was to shoot this match the day after the Colours match, ie. We would go to WTSC on Saturday and shoot the colours match with UCD, spend the night there, and then shoot the triad match the next day with UCD and WTSC. This was accepted as a good idea as it kept the social side of the colours match alive.

17 DIT Rifle Club

Mark reported on the meeting with UCD/WTSC/NTSA regarding air rifle, specifically the request from DIT for assistance in setting up a rifle club. They have no space at present for a rifle range and were wondering if we could let them use our range. This request provoked heated debate. Concerns were voiced regarding security; extra wear&tear on equipment; the additional amount of work DURC would have to do; and how college would react. Alan worried about the loss of the time on the range for the ROs and wondered what we would get in return for the effort. Ben pointed out that one person had to oversee them at all times. Deirdre wondered if the effort was worthwhile. John felt it would be nice to be able to help but was in agreement with the others. Mark expressed dissappointment with this general attitude and more heated debate followed. Ben wondered if we would be sacrificing the Pony Club efforts for DIT. Steven pointed

17 DIT Rifle Club

out that there was less politics involved in assisting the pony club. John pointed out that if a DIT member broke a DURC piece of equipment, the question of who pays to fix it would be politically tricky. Mark wondered if that could be agreed upon in advance. Steven said that unless we called in soliciters, that this would not be possible. Since the debate was running on, Steven called for a round-table expression of opinions. Alan thought it was a nice idea to help DIT but that it was unworkable. Deirdre agreed with him. Ben said that manpower wasn’t that big a problem, especially if UCDRC assist with range space. He felt we should not rule out letting them use our range space, but not to list it as a first option. Deirdre suggested a trial period and this was agreed as being a necessary step. John said that from the A&E officer’s viewpoint, they could not be allowed access to our safe or rifles as College authorities were allready uneasy with the idea of DURC, let alone another rifle club. He also felt they should be able to supply their own rifles and ammo. He suggested we assist as consultants instead of giving material support. Mark again expressed unhappiness with the general attitude towards the proposal, citing the assistance that DURC received upon it’s startup, the general view expressed over the years that more college clubs would be a good idea, the current support Geoff Cooney is showing towards the sport and he expressed his belief that most of the problems discussed assumed a lack of respect for our equipment from DIT. Steven said that we didn’t have a real manpower problem that prevented the proposal, that giving over range space was something he felt we could just about live with, but that we couldn’t let them use our rifles/equipment. He worried about the longevity of the DIT club and pointed out that so far we had not met these people, that only Denise O’Brien (UCD Captain) and Mark had. He felt that a half-way meeting might be better, that they get a little equipment and a little space and we help from there. John at this point asked to qualify his statement in light of Mark’s and said that he had no personal bias towards the people in the DIT club, but that normal people cause wear and tear on equipment and the situation with funding replacement gear was not a well-defined one. He was also worried that College would not permit this at all, and that should an accident happen, that while we have our 40-year history in the college to use as proof that we are normally safe, if DIT were to have an accident, that would not be the case. Ben pointed out that Denise had invited them to the meeting and thus UCDRC were keen on the idea, so why not use UCD for the range space and us to assist, so that both UCD and TCD meet DIT half-way instead of us shouldering the full burden.

At this point, Steven called for votes as the debate had now gone on for over thirty minutes. Do we want to allow DIT to use our range space

For  Against  Abstention
5  1  0

Motion carried

At this point there as some debate as to what this meant. It was clarified to mean, did we feel that DIT using our range was in our club’s best interests, and the votes were then checked and agreed on as before.

18 TTT

Do we want to let DIT use our rifles

For  Against  Abstention
0  6  0

Motion rejected Do we want to let DIT use our jackets?

For  Against  Abstention
0  6  0

Motion rejected

A suggestion from Steven was that we find the equipment cheaply for them and they raise the necessary funds. More discussion followed over our course of action and the following was agreed on:

DIT will attend both the WAIDAM Open and the DURC Open. Afterwards, they will shoot once in DURC and once in UCDRC. They can use our old jackets and range space, but not our rifles. After this, we will meet again to review the situation.

The above is contingent on finding out how much we have to charge them for range hire and who receives this money, as well as the relevant College authorities agreeing to the scheme. Ben pointed out that we must make all involved aware of the temporary nature of the scheme or College may think we don’t need all the time we have on the range at present.

18 TTT

Mark reported that the Trinity Top Ten competition was in dissarray. Few entries were received, and even fewer entry fees. Mark reported he was going to announce an amnesty period for competitiors to catch up with their cards and settle their entry fees, and that if at the end of this period matters were still unresolved that he would refund entry fees and cancel the competition. Steven suggested phoning around and checking on people’s status. Alan suggested just dropping the competition. Mark reminded everyone of the past problems running the TTT and said that if we dropped it this year, it was effectively over permenantly. It was discussed and the decision was that Mark would find out if EARC had paid their entry fees. If they had not, the TTT would be scrapped and entry fees repaid. Task: Find out if EARC will pay thier entry fee – if not, cancel TTT and repay entry fees. (Mark)

T.B.N.O.B.T.M.W.D.C.

Steven Watterson, Club Captain

Mark Dennehy, Club Secretary

18 TTT

Tasks agreed on at meeting